how many oxygen atoms are in 2.40 g of quartz?
Answer #1

Quartz formula is SiO2

no of oxygen atoms = mass of substance*6.023*1023 *atomicity/Gram molar mass

gram molar mass of SiO2 = 28+32 = 60g/mole

Atomicity of Oxygen in SiO2 = 2

no of oxygen atoms = mass of SiO2* 6.023*1023 * atomicity of oxygen in SiO2/Gram molar mass of SiO2

                               = 2.4*6.023*1023 *2/60   = 4.8*1022 oxygen atoms